UT 20130916 AO#2 DX

Observer:JCGuerra(LBTO),SEspostio,LBuzoni,RBruglio,APuglisi (Arcetri Observatory)
Telescope Operator:
SW Support:
AO Support:
Telescope Support:
Instrument Support: (LBTO)

Summary

  • Py-WFS set up an operations without any problem.
  • Installation of the RR without problem.
  • Check of the mouse hole mask ok.
  • Check of the reconstructor at bin 1 from yesterday.
  • Check pupil wobble with bin 1.
  • Analysis of the strehl data related to 30Hz vibration. Arcetri report.
  • Optimizing the AO configuration for Bin 2.
  • Acquisition of IRTC psf.
  • Reconstructor of 152 modes @ Bin2 ready to test on sky. Not yet
  • Pointing to star and not time to acuire star. Close dome due to clouds and humidity.
  • Hand over @ 5pm *
  • More details in the Arcetri report. Attached to this page

Detail

The FLAO system is power on and set up without any problem. The system is configured to work with Bin 1 and to test how it is working.

The bin 1 resconstructor is loaded and tested showing same performances and stability than yesterday, this is a good signal after the succersfull previous night on sky.

Board setup for daytime is saved; etro_ts4

The pupil wobble of the system was checked for bin 1 showing a good range of operations of the camara lens (CL).

Telescope Rot CL (x) CL(y)
300 0 82
210 32 30
120 17 -8
30 -50 57
-60 0 90
-40 0 100
-30 -23 93

A digital shift of one pixel is applied in the Y direction to adjust the pupil wobble and to get the maxiumum camara lens range

it has been done some analysis of data acquired yesterday to determine some relationship of the tree house (30Hz) frequency in depres the strehl ratio. Arcetri is preparating a more detailed report about it. IA new set of data it will be good to have it.

Bin2

Setup the system to work to acquire the bin 2 INTMAT.

The bin2 pupils are a trick process because it is not easy to match the actual subaperture and actuators symetry. for the bin 2 we have a rotation of about 60 degrees.

Several iterations it is done to acquire the bin 2 pupils. The first pupilwiht a reconstructor of 152 modes is giving problems in one the edges of the mirror, the loop breaks during the optimize gain. The pupil to test is 20130916_173526

A reconstructor is built; 20130916_174622

Test of the recosntructor with a disturbance of 0.8" and high flux.

TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20130916_180534 ??? 7.3 90 5 NaN -1.00 174622 2 153 991 0.6 0.7 0.5 3 2724 ON 57.6 H 40 250 ON NO  
20130916_180601 ??? 7.3 90 4 NaN -1.00 174622 2 153 991 0.6 0.7 0.5 3 2722 ON 56.7 H 40 250 ON NO  
20130916_180629 ??? 7.3 90 5 NaN -1.00 174622 2 153 991 0.6 0.7 0.5 3 2724 ON 55.0 H 40 250 ON NO  
20130916_180734 ??? 7.3 90 3 NaN -1.00 174622 2 153 991 0.6 0.7 0.5 3 2773 ON 55.8 H 40 250 ON NO  
20130916_180802 ??? 7.3 90 5 NaN -1.00 174622 2 153 991 0.6 0.7 0.5 3 2755 ON 57.6 H 40 250 ON NO  
20130916_180830 ??? 7.3 90 2 NaN -1.00 174622 2 153 991 0.6 0.7 0.5 3 2746 ON 56.7 H 40 250 ON NO  
20130916_181351 ??? 7.3 90 4 NaN -1.00 174622 2 153 991 0.6 0.7 0.5 3 2738 ON 55.3 H 40 250 ON NO  
20130916_181419 ??? 7.3 90 4 NaN -1.00 174622 2 153 991 0.6 0.7 0.5 3 2737 ON 56.6 H 40 250 ON NO  
20130916_181446 ??? 7.3 90 5 NaN -1.00 174622 2 153 991 0.6 0.7 0.5 3 2736 ON 59.0 H 40 250 ON NO  

With this pupils we check the pupil wobble.

Telescope Rot CL (x) CL(y)
30 19 31
-30 63 8
-60 82 -34
120 -12 -30
210 38 -56
300 76 -6

Check the performances of the reconstructor with the magnitude and frequency for the bin2. The flux is droped to match a 10 magnitude star. Three images were taken without disturbance. Applied disturbance with seeing 0.8"

TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
WARNING: DISTURB IS OFF!!
20130916_184442 ??? 7.3 90 4 NaN -1.00 174622 2 153 991 0.5 0.2 0.3 3 2850 ON 79.8 H 20 500 OFF NO  
WARNING: DISTURB IS OFF!!
20130916_184525 ??? 7.3 90 4 NaN -1.00 174622 2 153 991 0.5 0.2 0.3 3 2799 ON 77.6 H 20 500 OFF NO  
WARNING: DISTURB IS OFF!!
20130916_184607 ??? 7.3 90 5 NaN -1.00 174622 2 153 991 0.5 0.2 0.3 3 2786 ON 73.7 H 20 500 OFF NO  
20130916_185740 ??? 10.9 90 3 NaN -1.00 174622 2 153 991 0.7 0.5 0.4 3 103 ON 50.6 H 200 50 ON NO  
20130916_185756 ??? 10.9 90 4 NaN -1.00 174622 2 153 991 0.7 0.5 0.4 3 102 ON 48.1 H 200 50 ON NO  
20130916_185812 ??? 10.9 90 3 NaN -1.00 174622 2 153 991 0.7 0.5 0.4 3 101 ON 53.6 H 200 50 ON NO  
20130916_185854 ??? 10.9 90 3 NaN -1.00 174622 2 153 991 0.7 0.5 0.4 3 101 ON 52.2 H 200 50 ON NO  
20130916_185909 ??? 10.8 90 2 NaN -1.00 174622 2 153 991 0.7 0.5 0.4 3 104 ON 54.0 H 200 50 ON NO  
20130916_185925 ??? 10.8 90 3 NaN -1.00 174622 2 153 991 0.7 0.5 0.4 3 104 ON 47.3 H 200 50 ON NO  

Looking the PSD of the data we can see some oscillation at 60Hz and more data are collected to see this effect. This can be due to any event in the chamber, probably during the tomorrow close dome tests we can collect more data.

Track Number from:_190805 to _191315.

Closing the loop with a magnitude simulated of 11.7 and a disturbance of 0.8" seeing the CL went to end to the maximum range. this is not right and a pupil wobble was checked again. Several darks of the CCD39 were taken and a digital shift is applied to one pixel to the left. The new pupil is : _193614

Telescope Rot CL (x) CL(y)
100 -76 -16
160 -40 -80
60 -69 -36
0 -8 80

With the new pupil a new INTMAT and reconstructor is built.

The flux is dropped to simulated amgnitude to 11.7, the loop is close and we still see some effect in the edge of the mirror, To face again this problem one subaperture. The problem seems that it is still there.

Checking the pupils we have found that it has been a vignetting of the cube that obscure some subaperrture, so adjusting the cube and a new pupil was acquired. _220247

With the new pupil a new INTMAT and reconstructor were built;

Rec: _221257.

The reconstructoir is tested at low flux where we have found the problems. This reconstructor is working and it is not showing any problem ijn the edges.

mag 11 Ok mag 11.7 Ok

TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20130916_222507 ??? 11.9 90 3 NaN -1.00 221257 2 153 625 0.5 0.6 0.4 3 62 ON 35.7 H 300 30 ON NO  
20130916_222522 ??? 12.0 90 7 NaN -1.00 221257 2 153 625 0.5 0.6 0.4 3 61 ON 36.4 H 300 30 ON NO  
20130916_222537 ??? 12.0 90 4 NaN -1.00 221257 2 153 625 0.5 0.6 0.4 3 60 ON 35.1 H 300 30 ON NO  
20130916_222616 ??? 12.8 90 4 NaN -1.00 221257 2 153 625 0.5 0.6 0.4 3 29 ON 38.4 H 300 30 ON NO - Pupils not on target
20130916_222901 ??? 12.0 90 5 NaN -1.00 221257 2 153 625 0.5 0.6 0.4 3 61 ON 37.7 H 300 30 ON NO  
20130916_222916 ??? 12.0 90 5 NaN -1.00 221257 2 153 625 0.5 0.6 0.4 3 61 ON 40.3 H 300 30 ON NO  
20130916_222931 ??? 12.0 90 7 NaN -1.00 221257 2 153 625 0.5 0.6 0.4 3 61 ON 39.1 H 300 30 ON NO  

Go to high flux and 1kHz to estimate better the performances a disturbance is applied of 0.8" seeing

The Reconstructor is ready to be tested on sky.

On sky

Telescope point to a 10.7 magnitude star. At the moment ot acquire the star, we have to close due to clouds and humidity..... Hoope tomorrow a better luck..

--++AO886 10.7 magnitdue

Orientation of telescope pointing and CCD47

Up IE down CCD47

right/left not yet defined

-- JuanGuerra - 17 Sep 2013
  • bin2 reconstructor good to test on sky:
    221257_bin2_0.png
I Attachment Action Size Date Who Comment
221257_bin2_0.pngpng 221257_bin2_0.png manage 51 K 17 Sep 2013 - 06:41 UnknownUser bin2 reconstructor good to test on sky
Topic revision: r3 - 21 Sep 2013, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback